I have been trying to solve this question but I cant figure it out: The sum of the present ages of Eric and his father is 58 years. In 10 years, his father will be twice as old as Eric will be at that time. Find their present ages.
Answer by stanbon(75887)

I have been trying to solve this question but I cant figure it out: The sum of the present ages of Eric and his father is 58 years. In 10 years, his father will be twice as old as Eric will be at that time. Find their present ages.
------------------------
EQUATIONS:
Present: F + E = 58
Future: (F+10) = 2(E+10)
------------------------
Rearrange:
F + E = 58
F -2E = 10
-------------
Subtract the 2nd equation from the 1st to solve for E:
3E = 48
E = 16 yrs old (Eric's age now)
Since F+E= 58, F=42 (Father's age now)
========================
Cheers,
Stan H.